OpenStudy (anonymous):

1/6x^2 = 1/3x^2 - 1/x

OpenStudy (anonymous):

so you wanna find value of x

OpenStudy (anonymous):

\[\frac{ 1 }{ 6x^2 } =\frac{ 1 }{ 3x^2 } - \frac{ 1 }{ x }\]

OpenStudy (anonymous):

hmm so you tried simplifying? like if u cancel x from denominator you will get 1/6x =1/3x -1

OpenStudy (anonymous):

looks easier now ,doesnt it? take 1/3x on other side take 1/3x common you will get 1/3x(1/2 -1) = -1

OpenStudy (anonymous):

can you solve the rest?

OpenStudy (anonymous):

yes thanks!

OpenStudy (anonymous):

you will get -1/6x = -1 1/6x=1 6x=1 so x=1/6

OpenStudy (anonymous):

okay good :)